OpenStudy (anonymous):

11z/16 + 7z/8 = 5/16

OpenStudy (anonymous):

@cj49 @Teddyiswatshecallsme

OpenStudy (cj49):

11z/16+7z/8 11z+14z/16=5/16 25z/16=5/16 25z=5 z=1/5

OpenStudy (teddyiswatshecallsme):

What are you trying to do here?